OpenStudy (anonymous):

Simplify the expression 2+3i/4+3i

OpenStudy (anonymous):

\[\frac{ 2+3i }{ 4+3i }*\frac{ 4-3i }{ 4-3i }\] \[=\frac{ 8+6i-9i^2 }{ 16-9i^2 }\] i^2=-1 \[=\frac{ 8+6i+9 }{ 16+9 }\] \[=\frac{ 17+6i }{ 25 }\]

OpenStudy (anonymous):

http://www.wolframalpha.com/